# Hybrid classical-quantum communication networks

### v1

@@ 2.1.5 All-optical signal processing
Optical TDM, SDM, and the various flavors of WDM described in the previous section form a subset of a wider corpus of techniques known collectively as all-optical signal processing [251]. Broadly defined, all-optical signal processing encompasses any communication operation—e.g., multiplexing, impairment compensation, and logic—that is performed entirely in the optical regime with no conversion either to or from the electronic domain. Significant experimental progress has been made on a variety of photonic capabilities conventionally associated with electronics, including demultiplexing and pulse shaping in more complex multiuser paradigms such as optical code-division multiple-access [107], optical orthogonal frequency-division multiplexing [105], and optical orthogonal time-division multiplexing [203]; wavelength exchange and multicasting [20, 253, 31, 141]; tunable delay lines [169, 220]; and high-speed logic [256, 75].
@@
With the exception of WDM systems and the continued trend to supplant electrical wires with optical fiber for data transmission on ever-shortening length scales, such as datacenters and supercomputers [223, 155, 38], all-optical signal processing techniques have faced similar challenges toward adoption as optical packet switching. And in fact, in modern lightwave communications [13]—particularly in the explosion of coherent encoding formats [119]—digital signal processing (DSP) has improved to such an extent that the trend has even gone in the opposite direction, with electronics assuming responsibilities previously the purview of photonics. For example, rather than compensating dispersive spreading with optical components such as dispersion-compensating fibers and chirped fiber Bragg gratings, chromatic dispersion is now routinely compensated digitally after OE conversion. More generally, the expansion of forward-error correction [47] has massively relaxed SNR requirements in optical transmission, shifting much of the burden for error-free communications into the digital domain. Interestingly the destruction of quantum coherence inherent in photodetection precludes the application of most of these DSP techniques for quantum signals, inspiring us to speculate that quantum networks might finally provide the application needed to push all-optical processing from the lab to the field.
@@ 2.2.1 Basic concepts
Classical optical communication protocols are typically classified based on their detection methods: direct detection and coherent detection. In contrast, quantum communication protocols are generally divided into two categories based on the dimensionality of the information carriers: discrete-variable (DV) and continuous-variable (CV) protocols. DV protocols utilize information carriers defined in finite-dimensional spaces, such as the polarization or time-bin of single photons, whereas CV protocols rely on information carriers defined in infinite-dimensional spaces, like the quadrature components of an optical mode.
@@
Light Sources Laser; Incoherent light source Laser; Incoherent light source; Single-photon source; Squeezed state of light; Entangled light
Information Encoding Actively encodes classical information on orthogonal states of light. Actively encodes classical or quantum information on orthogonal or non-orthogonal states of light; Passively generates information during measurement process (e.g., entanglement-based QKD)
Photon Detection Phase-insensitive intensity detection; Phase-sensitive coherent detection Phase-insensitive single-photon detection; Phase-sensitive coherent detection
@@ 2.2.2 Quantum repeaters
The idea of that something other than phase-insensitive amplification is required for large-scale high-performance quantum communications is a direct consequence of the no-cloning theorem[254, 60]. However, it is useful to consider the need in the context of an application in a concrete use case. That use case was the protocol of QKD, which is covered subsequently, where a great deal of work went into developing better protocols and systems to transmit higher rates of keys over increasingly larger-loss channels. A recurring output was that the rate decreased exponentially with the increasing channel length. It was some time before it was established that a quantum repeater, or something similar, was rigorously needed to extend the performance of quantum key exchange, even with prefect hardware due to a fundamental rate-loss trade-off [228]. Later, this result was generalized more broadly for quantum communications, further cementing the need for a quantum repeater to realize long-distance quantum communications[179].
@@
As of the time of writing, there have been more than a dozen related review papers summarized by a recent review of quantum repeaters[8]. The first quantum repeater concepts are based on entanglement distribution, purification and quantum teleportation[27, 56]. Here quantum communications can be achieved by distributing entanglement and performing quantum teleportation[23] with the distributed shared entanglement. To deal with the nondeterministic throughput due to transmission loss, quantum memory is used to hold successfully distributed entangled resources until they are needed for quantum teleportation. To reduce the impact of exponentially increasing fiber loss, the channel is broken into smaller, lower-loss links. To mitigate the decoherence of quantum storage, entanglement purification protocols can transform a number of imperfect input entangled resources into a fewer number of higher quality entangled resources for quantum teleportation.
@@
To capture the major differences in the development of quantum repeater (QR) protocols, three generations have been identified by Munro and colleagues[152] as well as by Muralidharan and colleagues[156]. As characterized in Figure 4, the different generations of QRs may be thought of by how they handle loss and operational errors, as well as the required classical communications (i.e., one-way or two-way signaling). For first and second generation QRs, loss errors are corrected via Heralded Entanglement Generation (HEG), while third generation QRs utilize Quantum Error Correction (QEC). For operational errors, Heralded Entanglement Purification (HEP) is the exclusive domain of first generation QRs, while second and third generations use QEC.
@@
While the development of QRs is a very active area of research, their integration into classical networks is likely to be a strong function of the underlying approach. For example, some challenges first identified in the context of “coexistence” of classical and quantum channels for QKD, such as Raman scattering, may drive some QR concepts towards using dedicated dark fiber for the quantum transmission while others, may coexist. Though it seems likely that some classical level signals for quantum channel control and stabilization will be needed even if it is advantageous to refrain from using remaining bandwidth to send additional classical data.
@@
Figure 4: Quantum repeater protocols by generation. Image reproduced under a Creative Commons Attribution 4.0 International License (http://creativecommons.org/licenses/by/4.0/) from [156]. Key steps include Heralded Entanglement Generation (HEG), Heralded Entanglement Purification (HEP), and Quantum Error Correction (QEC).
@@ 2.2.3 Near-term application—quantum key distribution
The first QKD protocol, the celebrated BB84 QKD protocol  [21], was proposed by Bennett and Brassard at a time when it was assumed that the key distribution problem in cryptography was already solved by using public-key cryptosystems, such as the RSA algorithm  [201]. Potential real-life applications of QKD were not recognized until quantum researchers realized that quantum computing might pose a formidable challenge to contemporary public cryptographic protocols, including RSA, as evidenced by Shor’s algorithm for factoring  [213]. This spurred the rapid growth of QKD and related quantum cryptographic protocols, such as quantum secret sharing (QSS)  [90, 40, 121], quantum digital signatures (QDS)  [82, 66], etc. Today, commercial QKD systems are readily available, and several large-scale QKD networks have been deployed worldwide  [73, 181, 43, 209, 139, 69].
@@
In classical optical communication, the two most commonly used detection schemes are direct detection and coherent detection. Their counterparts in QKD are single-photon detection and optical coherent detection, respectively. QKD protocols are generally categorized into two families: DV QKD and CV QKD [259], based on the dimensionality of the encoding space. However, the defining characteristics of different QKD protocols are determined primarily by the detection scheme used.
Single-photon detectors are commonly employed to implement DV-QKD, where the dimension of the encoding space (such as the polarization of a single photon) is finite, although they have also been applied in certain CV-QKD protocols, such as time-frequency coding QKD [185, 186, 168, 165], where the dimension of the encoding space is infinite. In contrast, optical coherent detectors are almost exclusively used in CV-QKD protocols based on quadrature encoding using coherent states or squeezed states [193, 97, 88]. An interesting variation is operating a coherent receiver in phase-insensitive photon detection mode [191] to implement DV-QKD protocols  [188]. Such an approach may inherit certain advantages of both detection schemes.
@@
QKD is fundamentally a point-to-point protocol. To fully realize its potential, large-scale QKD networks with broad geographical coverage, accessible to a large number of users, are required. However, in the absence of QRs, there are inherent limitations on the achievable quantum communication rates over a lossy bosonic channel. This understanding emerged from the historical effort to determine the secure key rate achievable using existing QKD protocols. The fundamental rate-loss tradeoff was first established by Takeoka, Guha, and Wilde in the context of secret key agreement over a noisy lossy channel, leading to the “TGW” bound [228]. Subsequently, Pirandola, Laurenza, Ottaviani, and Banchi expanded the analysis to cover a broader range of quantum channels, culminating in the ’PLOB’ bound for quantum repeaterless channels [179].
@@
Similar to classical networking, an efficient many-user network in quantum communication depends on a well-designed switching (or routing) scheme. Most existing QKD networks are structured like “circuit-switched” networks in classical communication: before two remote parties can begin QKD, a direct optical link must be established and reserved for the duration of the quantum communication phase. More recently, the packet-switching scheme, which has been shown to offer better efficiency in classical networking, has been introduced into quantum communication and QKD [62, 153].
@@
A critical component of packet-switched quantum networks is the quantum counterpart of the reconfigurable optical add-drop multiplexer (ROADM), known as the q-ROADM, illustrated in Fig. 5. Although it is too early to present a detailed design, developing a general-purpose q-ROADM will demand significant advancements in fundamental quantum technologies, including long-lived quantum memory and quantum repeaters.
Figure 5: A conceptual design of a quantum ROADM. The classical header can be measured and regenerated in the classical processing unit, while the quantum payload is stored, corrected for errors, and, if needed, undergoes frequency conversion in the quantum processing unit. (Images reproduced with permission from a Creative Commons Attribution 4.0 International License (http://creativecommons.org/licenses/by/4.0/) [62])
@@ 2.2.4 Quantum communication networks
In view of such limitations, multiplexing strategies based on classical lightwave communications [cf. Sec. 2.1.4] have been increasingly adopted in the quantum domain. As general guidelines, any prospective “quantum-capable” multiplexing approach should be (i) all optical in the broad sense (i.e., implementable without OE conversion) and (ii) avoid amplification (i.e., amplification must not be intrinsic to its operation); standard OE conversion represents a measurement that collapses the quantum state, and qubits cannot be amplified without introducing noise, a consequence of the no-cloning theorem [254]. Thankfully, WDM does meet these requirements and provides a powerful toolkit for allocating quantum resources.
@@
In particular, the frequency entanglement generated through narrowband-pumped spontaneous parametric downconversion [212] offers broadband spectral correlations that serve as natural resources for entangling many users; for example, when combined with entanglement in another DoF such as polarization or time bins, pairs of frequency-entangled bands can be allocated to any two parties wishing to share entanglement. This approach was demonstrated in a fully connected four-user QKD experiment facilitated by a shared polarization-frequency hyperentangled photon source and a network of interconnected DWDM filters: first to demultiplex twelve 100 GHz-wide bands, and then to multiplex channel triplets destined for each receiver [249]. In analogy with classical WDM, this experiment can be viewed as a realization of “fixed DWDM,” where the allocation is static and each channel occupies a uniform spectral slot.
@@
More recent experiments have expanded on this concept by bringing elastic optical networking concepts squarely into the quantum domain; replacing fixed DWDMs with a single WSS was shown to enable multiuser entanglement distribution in which each pair of users is assigned bandwidth dynamically and in proportion to demand [138, 2]. Although WSSs impart higher losses ($\sim$5 dB) compared to a single DWDM ($\sim$0.5 dB), they can support nodes without additional components, thus offering better scaling in the high-user regime [138]. For example, a recent experiment merging a broadband Sagnac polarization-entangled source [16] with C- and L-band WSSs demonstrated high-fidelity entanglement in 150 pairs of 25 GHz-wide channels, which when combined with the high port counts (e.g., 20 and beyond) and fine spectral granularity (e.g., 3.125 GHz) available from modern WSSs [159] point to a promising future for flex-grid quantum networking.
@@
In such a future, one can envision leveraging multiple WSSs connected together to support complex mesh or multihop network topologies. Fig. 6(a) highlights one simple example, whereby a source of frequency-entangled photons distributed by a single WSS in a star configuration (top) is expanded to service a second set of nodes of with a second WSS (bottom). Assuming these WSSs are associated with distinct QLANs, for example, this construction represents perhaps the simplest quantum internetwork—i.e., a true “quantum Internet” [19]. Depending on the design of the optical source, quantum information can be carried in any DoF compatible with the fiber transmission medium; Fig. 6(b) depicts a version in which three separate pairs of polarization-frequency hyperentangled bands (red, green, and blue) are allocated to supply polarization entanglement to all pairs of users in a three-node network [1].
@@
More broadly, WSSs form just one example—albeit a particularly sophisticated one—of a variety of spatial switching technologies available for quantum networking, including microelectromechanical and optomechanical versions as well [234, 184, 51, 92, 142]. In all cases, the reconfigurability available permits dynamic allocation of bandwidth to pairs of users on demand, facilitating any-to-any connectivity without the dedicated wavelengths and extra background light inherent to a fully connected quantum network in the fixed DWDM paradigm [19, 1].
@@
Nonetheless, it appears to us that the fundamental requirements of quantum networking furnish the strongest motivation to date for developing all-optical signal processing approaches. Indeed, both quantum frequency conversion and quantum transduction [126, 143, 3, 95] will be critical for interfacing stationary quantum systems with telecom-band photons, required irrespective of whether circuit- or packet-switching is employed. And an extensive toolkit for quantum gates in SMF-compatible DoFs like time bins [100, 227], frequency bins [123, 146, 176, 140, 134], and orthogonal modes [35, 7, 202, 112]—as well as MMF-compatible DoFs like transverse spatial modes [58, 239]—has emerged in recent years. Although experiments so far have demonstrated local quantum information processing functionalities rather than distributed networking per se, the nonlinear, electro-optic, and pulse shaping techniques employed have significant potential in this space; for example, the frequency-bin discrete Fourier transforms and interleaved beamsplitter—both recently experimentally realized via the electro-optic-based quantum frequency processor (QFP) [136, 135]—form potential building blocks for spectrally multiplexed quantum interconnects [134]. The degree to which these spectro-temporal and spatial DoFs will assist in not only multiplexing quantum information but also in encoding the information itself remains an open question and fertile ground for continued research.

### v2

@@ 2.1.5 All-optical signal processing
Optical TDM, SDM, and the various flavors of WDM described in Sec. 2.1.4 form a subset of a wider corpus of techniques known collectively as all-optical signal processing [54]. Broadly defined, all-optical signal processing encompasses any communication operation—e.g., multiplexing, impairment compensation, and logic—that is performed entirely in the optical regime with no conversion either to or from the electronic domain. Significant experimental progress has been made on a variety of photonic capabilities conventionally associated with electronics, including demultiplexing and pulse shaping in more complex multiuser paradigms such as optical code-division multiple-access [55], optical orthogonal frequency-division multiplexing [56], and optical orthogonal time-division multiplexing [57]; wavelength exchange and multicasting [58, 59, 60, 61]; tunable delay lines [62, 63]; and high-speed logic [64, 65].
@@
With the exception of WDM systems and the continued trend to supplant electrical wires with optical fiber for data transmission on ever-shortening length scales, such as datacenters and supercomputers [72, 73, 74], all-optical signal processing techniques have faced similar challenges toward adoption as optical packet switching. And in fact, in modern lightwave communications [6]—particularly in the explosion of coherent encoding formats [19]—DSP has improved to such an extent that the trend has even gone in the opposite direction, with electronics assuming responsibilities previously the purview of photonics. For example, rather than compensating dispersive spreading with optical components such as dispersion-compensating fibers and chirped fiber Bragg gratings, chromatic dispersion is now routinely compensated digitally after OE conversion. More generally, the expansion of forward-error correction [75] has massively relaxed SNR requirements in optical transmission, shifting much of the burden for error-free communications into the digital domain. Interestingly, the destruction of quantum coherence inherent in photodetection precludes the application of most of these DSP techniques for quantum signals, inspiring us to speculate that quantum networks might finally provide the application needed to push all-optical processing from the lab to the field.
@@ 2.2.1 Basic concepts
Classical optical communication protocols are typically classified based on their detection methods: direct detection and coherent detection. In contrast, quantum communication protocols are generally divided into two categories based on the dimensionality of the information carriers: discrete-variable (DV) and continuous-variable (CV) protocols. DV protocols utilize information carriers defined in finite-dimensional spaces, such as the polarization or time bin of single photons, whereas CV protocols rely on information carriers defined in infinite-dimensional spaces, like the quadrature components of an optical mode.
@@
Light Source Laser; incoherent light source. Laser; incoherent light source; single-photon source; squeezed state of light; entangled light.
Information Encoding Actively encodes classical information on orthogonal states of light. Actively encodes classical or quantum information on orthogonal or non-orthogonal states of light; passively generates information during measurement process (e.g., entanglement-based QKD).
Photon Detection Phase-insensitive intensity detection; phase-sensitive coherent detection. Phase-insensitive single-photon detection; phase-sensitive coherent detection.
@@ 2.2.2 Quantum repeaters
The idea of that something other than phase-insensitive amplification is required for large-scale high-performance quantum communications is a direct consequence of the no-cloning theorem [76, 77]. However, the need developed historically in the context of a concrete use case. That use case was the protocol of QKD, which is covered subsequently, where a great deal of work went into developing better protocols and systems to transmit higher rates of keys over increasingly larger-loss channels. A recurring output was that the rate decreased exponentially with the increasing channel length. It was some time before it was established that a quantum repeater, or something similar, was rigorously needed to extend the performance of quantum key exchange, even with prefect hardware due to a fundamental rate-loss tradeoff [83]. Later, this result was generalized more broadly for quantum communications, further cementing the need for a quantum repeater to realize long-distance quantum communications[84].
As of the time of writing, there have been more than a dozen related review papers summarized by a recent review of quantum repeaters [85].
@@
The first quantum repeater concepts are based on entanglement distribution, purification and quantum teleportation [86, 87]. Here quantum communications can be achieved by distributing entanglement and performing quantum teleportation [81] with the distributed shared entanglement. To deal with the nondeterministic throughput due to transmission loss, quantum memory is used to hold successfully distributed entangled resources until they are needed for quantum teleportation. To reduce the impact of exponentially increasing fiber loss, the channel is broken into smaller, lower-loss links.
To mitigate the decoherence of quantum storage, entanglement purification protocols can transform a number of imperfect input entangled resources into a fewer number of higher quality entangled resources for quantum teleportation.
@@
To capture the major differences in the development of quantum repeater (QR) protocols, three generations have been identified by Munro and colleagues [88] as well as by Muralidharan and colleagues [89]. As characterized in Fig. 4, the different generations of QRs may be thought of by how they handle loss and operational errors, as well as the required classical communications (i.e., one-way or two-way signaling). For first and second generation QRs, loss errors are corrected via heralded entanglement generation (HEG), while third generation QRs utilize quantum error correction (QEC). For operational errors, heralded entanglement purification (HEP) is the exclusive domain of first generation QRs, while second and third generations use QEC.
Experimental implementations of quantum repeaters still lag behind theoretical advances. To date, most demonstrations have been limited to entangling a small number of distant quantum memories—a key milestone toward realizing first-generation quantum repeaters [90, 91, 92, 93]. While scalable, practical quantum repeaters remain out of reach, the technologies developed so far have already shown promise for near-term applications, such as device-independent quantum key distribution (DI-QKD) [94], which can establish security without requiring trust in most components of the QKD system. Furthermore, it has been shown that even a single repeater node can enhance communication rates: for instance, the secret key rate in QKD can surpass the fundamental limit of repeaterless protocols, improving from a linear $\eta$ scaling to a more favorable $\sqrt{\eta}$ scaling, where $\eta$ denotes the channel transmittance [95, 96].
@@
While the development of QRs is a very active area of research, their integration into classical networks is likely to be a strong function of the underlying approach. For example, some challenges first identified in the context of “coexistence” of classical and quantum channels for QKD, such as Raman scattering, may drive some QR concepts towards using dedicated dark fiber for the quantum transmission, while others may coexist. Either way, it seems likely that some classical-level signals for quantum channel control and stabilization will be needed even if it is advantageous to refrain from using remaining bandwidth to send additional classical data.
@@
Figure 4: Quantum repeater protocols by generation. Image reproduced from [89] under a Creative Commons Attribution 4.0 International License (http://creativecommons.org/licenses/by/4.0/). Key steps include heralded entanglement generation (HEG), heralded entanglement purification (HEP), and quantum error correction (QEC).
@@ 2.2.3 Near-term application—quantum key distribution
The first QKD protocol, the celebrated BB84 [10], was proposed by Bennett and Brassard at a time when it was assumed that the key distribution problem in cryptography was already solved by using public-key cryptosystems, such as the RSA algorithm  [98]. Potential real-life applications of QKD were not recognized until quantum researchers realized that quantum computing might pose a formidable challenge to contemporary public cryptographic protocols, including RSA, as evidenced by Shor’s algorithm for factoring  [99]. This spurred the rapid growth of QKD and related quantum cryptographic protocols, such as quantum secret sharing (QSS)  [100, 101, 102], quantum digital signatures (QDS)  [103, 104], etc. Today, commercial QKD systems are readily available, and several large-scale QKD networks have been deployed worldwide  [105, 106, 107, 108, 109, 110].
@@
In classical optical communication, the two most commonly used detection schemes are direct detection and coherent detection (cf. Fig. 3). Their counterparts in QKD are single-photon detection and optical coherent detection, respectively. QKD protocols are generally categorized into two families: DV-QKD and CV-QKD [97], based on the dimensionality of the encoding space. However, the defining characteristics of different QKD protocols are determined primarily by the detection scheme used. Table 2 summarizes the main features of several practical QKD protocols. For more details on practical implementations of QKD and QKD networks, readers are referred to recent review papers [97, 115].
Table 2: Some Practical QKD Protocols
Protocols Main Features
Decoy-State BB84 QKD [116, 117, 118] (a) The decoy-state scheme enables the implementation of single-photon-based protocols using conventional laser sources without significantly compromising performance. Decoy-state BB84 QKD is one of the most extensively studied and widely deployed DV-QKD protocols. (b) A secret key rate of 13.72 Mb/s has been demonstrated over a 2 dB channel using InGaAs avalanche photodiode detectors [119], and a key rate of 115.8 Mb/s over 10 km of standard optical fiber has been achieved using multipixel superconducting nanowire single-photon detectors [120]. (c) Using a three-state time-bin protocol combined with a one-decoy approach, secure QKD over a distance of 421 km has been demonstrated using ultralow-loss fiber [121]. Remarkably, satellite-to-ground QKD has also been demonstrated over a 1200km free-space channel [122].
Gaussian-Modulated Coherent States (GMCS) QKD [123] (a) The Gaussian-modulated coherent-state (GMCS) QKD protocol, based on conventional lasers, modulators, and optical homodyne detectors, is one of the most extensively studied CV-QKD schemes and is considered a cost-effective solution for practical deployment. (b) The mode-selection function of the LO in coherent detection makes GMCS QKD particularly attractive for classical and quantum co-existence networks [20]. (c) The room-temperature operation of optical homodyne detectors paves the way for on-chip QKD implementations [124], especially when the LO is generated at the receiver’s end [125, 126]. (d) Without detector deadtime, GMCS QKD can achieve very high secret key rates over low-loss channels. Recently, using a passive-state-preparation scheme [127], the authors in [128] estimated a gigabit-per-second secret key rate could be achieved over short distances. (e) Due to the unavoidable presence of vacuum noise, GMCS QKD typically operates in a very low SNR regime, making it less suitable for channels with high loss. The current maximum distance achieved over ultralow-loss fiber is 202 km [129].
Measurement-device-independent (MDI) QKD [130] (a) Measurement-device-independent (MDI) QKD enables secret key distribution without requiring trust in the measurement device, rendering it inherently immune to all detector side-channel attacks. (b) MDI-QKD networks also allow multiple users to share a centralized, high-cost detector, and have been successfully demonstrated over deployed optical fiber networks [131]. (c) Several variants of MDI-QKD, including twin-field QKD [132] and mode-pairing QKD [133, 134], improve the key rate scaling from linear with channel transmittance $\eta$ to a square-root dependence $\sqrt{\eta}$, leading to a recent demonstration of QKD over 1000 km of ultralow-loss fiber [135].
Single-photon detectors are commonly employed to implement DV-QKD, where the dimension of the encoding space (such as the polarization of a single photon) is finite, although they have also been applied in certain CV-QKD protocols, such as time-frequency-coded QKD [136, 137, 138, 139], where the dimension of the encoding space is infinite. In contrast, optical coherent detectors are almost exclusively used in CV-QKD protocols based on quadrature encoding using coherent states or squeezed states [140, 141, 123]. An interesting variation is operating a coherent receiver in phase-insensitive photon detection mode [142] to implement DV-QKD protocols [143]. Such an approach may inherit certain advantages of both detection schemes.
@@
QKD is fundamentally a point-to-point protocol. To fully realize its potential, large-scale QKD networks with broad geographical coverage, accessible to a large number of users, are required. However, in the absence of QRs, there are inherent limitations on the achievable quantum communication rates over a lossy bosonic channel. This understanding emerged from the historical effort to determine the secure key rate achievable using existing QKD protocols. The fundamental rate-loss tradeoff was first established by Takeoka, Guha, and Wilde in the context of secret key agreement over a noisy lossy channel, leading to the “TGW” bound [83]. Subsequently, Pirandola, Laurenza, Ottaviani, and Banchi expanded the analysis to cover a broader range of quantum channels, culminating in the “PLOB” bound for quantum repeaterless channels [84].
@@
Similar to classical networking, an efficient many-user network in quantum communication depends on a well-designed switching (or routing) scheme. Most existing QKD networks are structured like “circuit-switched” networks in classical communication: before two remote parties can begin QKD, a direct optical link must be established and reserved for the duration of the quantum communication phase. More recently, packet switching, which has been shown to offer better efficiency in classical networking, has been introduced into quantum communication and QKD [150, 151].
@@
A critical component of packet-switched quantum networks is the quantum counterpart of the reconfigurable optical add-drop multiplexer (ROADM), known as the q-ROADM, illustrated in Fig. 5. Although it is too early to present a detailed design, developing a general-purpose q-ROADM will demand significant advancements in fundamental quantum technologies, including long-lived quantum memories and QRs.
Figure 5: A conceptual design of a quantum ROADM. The classical header can be measured and regenerated in the classical processing unit, while the quantum payload is stored, corrected for errors, and, if needed, undergoes frequency conversion in the quantum processing unit. Image reproduced with permission from a Creative Commons Attribution 4.0 International License (http://creativecommons.org/licenses/by/4.0/[150].
@@ 2.2.4 Quantum communication networks
In view of such limitations, multiplexing strategies based on classical lightwave communications (cf. Sec. 2.1.4) have been increasingly adopted in the quantum domain. As general guidelines, any prospective “quantum-capable” multiplexing approach should be (i) all-optical in the broad sense (i.e., implementable without OE conversion) and (ii) avoid amplification (i.e., amplification must not be intrinsic to its operation). Standard OE conversion represents a measurement that collapses the quantum state, and qubits cannot be amplified without introducing noise [76, 77]. Thankfully, WDM does meet these requirements and provides a powerful toolkit for allocating quantum resources.
@@
In particular, the frequency entanglement generated through narrowband-pumped spontaneous parametric downconversion [153] offers broadband spectral correlations that serve as natural resources for entangling many users; for example, when combined with entanglement in another DoF such as polarization or time bins, pairs of frequency-entangled bands can be allocated to any two parties wishing to share entanglement. This approach was demonstrated in a fully connected four-user QKD experiment facilitated by a shared polarization—frequency hyperentangled photon source and a network of interconnected DWDM filters: first to demultiplex twelve 100 GHz-wide bands, and then to multiplex channel triplets destined for each receiver [154]. In analogy with classical WDM, this experiment can be viewed as a realization of “fixed DWDM,” where the allocation is static and each channel occupies a uniform spectral slot.
@@
More recent experiments have expanded on this concept by bringing elastic optical networking concepts squarely into the quantum domain; replacing fixed DWDMs with a single WSS was shown to enable multiuser entanglement distribution in which each pair of users is assigned bandwidth dynamically and in proportion to demand [155, 156]. Although WSSs impart higher losses ($\sim$5 dB) compared to a single DWDM ($\sim$0.5 dB), they can support nodes without additional components, thus offering better scaling in the high-user regime [155]. For example, a recent experiment merging a broadband Sagnac polarization-entangled source [157] with C- and L-band WSSs demonstrated high-fidelity entanglement in 150 pairs of 25 GHz-wide channels, which when combined with the high port counts (e.g., $\geq$20) and fine spectral granularity (e.g., 3.125 GHz) available from modern WSSs [48] point to a promising future for flex-grid quantum networking.
@@
In such a future, one can envision leveraging multiple WSSs connected together to support complex mesh or multihop network topologies. Figure 6(a) highlights one simple example, whereby a source of frequency-entangled photons distributed by a single WSS in a star configuration (top) is expanded to service a second set of nodes of with a second WSS (bottom). Assuming these WSSs are associated with distinct quantum local area networks (QLANs), for example, this construction represents perhaps the simplest quantum internetwork—i.e., a true “quantum Internet” [158]. Depending on the design of the optical source, quantum information can be carried in any DoF compatible with the fiber transmission medium; Fig. 6(b) depicts a version in which three separate pairs of polarization-frequency hyperentangled bands (red, green, and blue) are allocated to supply polarization entanglement to all pairs of users in a three-node network [159].
@@
More broadly, WSSs form just one example—albeit a particularly sophisticated one—of a variety of spatial switching technologies available for quantum networking, including microelectromechanical and optomechanical versions [160, 161, 162, 163, 164]. In all cases, the reconfigurability available permits dynamic allocation of bandwidth to pairs of users on demand, facilitating any-to-any connectivity without the dedicated wavelengths and extra background light inherent to a fully connected quantum network in the fixed DWDM paradigm [158, 159].
@@
Nonetheless, it appears to us that the fundamental requirements of quantum networking furnish the strongest motivation to date for developing all-optical signal processing approaches. Indeed, both quantum frequency conversion and quantum transduction [178, 179, 180, 181] will be critical for interfacing stationary quantum systems with telecom-band photons, required irrespective of whether circuit or packet switching is employed. And an extensive toolkit for quantum gates in SMF-compatible DoFs like time bins [182, 183], frequency bins [184, 185, 186, 187, 188], and orthogonal modes [189, 190, 191, 192]—as well as MMF-compatible DoFs like transverse spatial modes [193, 194]—has emerged in recent years. Although experiments so far have demonstrated local quantum information processing functionalities rather than distributed networking per se, the nonlinear, electro-optic, and pulse shaping techniques employed have significant potential in this space; for example, the frequency-bin discrete Fourier transform and interleaved beamsplitter—both recently experimentally realized via the electro-optic-based quantum frequency processor (QFP) [195, 196]—form potential building blocks for spectrally multiplexed quantum interconnects [188]. The degree to which these spectro-temporal and spatial DoFs will assist not only in multiplexing quantum information but also in encoding the information itself remains an open question and fertile ground for continued research.